### v2

# PackUV: Packed Gaussian UV Maps for 4D Volumetric Video

Abstract
Volumetric videos offer immersive 4D experiences, but remain difficult to reconstruct, store, and stream at scale. Existing Gaussian Splatting based methods achieve high-quality reconstruction but break down on long sequences, temporal inconsistency, and fail under large motions and disocclusions. Moreover, their outputs are typically incompatible with conventional video coding pipelines, preventing practical applications. We introduce PackUV, a novel 4D Gaussian representation that maps all Gaussian attributes into a sequence of structured, multi-scale UV atlases, enabling compact, image-native storage. To fit this representation from multi-view videos, we propose PackUV-GS, a temporally consistent fitting method that directly optimizes Gaussian parameters in the UV domain. A flow-guided Gaussian labeling and video keyframing module identifies dynamic Gaussians, stabilizes static regions, and preserves temporal coherence even under large motions and disocclusions. The resulting UV atlas format is the first unified volumetric video representation compatible with standard video codecs (e.g., FFV1) without losing quality, enabling efficient streaming within existing multimedia infrastructure. To evaluate long-duration volumetric capture, we present PackUV-2B, the largest multi-view 4D dataset to date, featuring more than 50 synchronized cameras, substantial motion, and frequent disocclusions across 100 sequences and 2B (billion) frames. Extensive experiments demonstrate that our method surpasses existing baselines in rendering fidelity while scaling to sequences up to 30 minutes with consistent quality.
Figure 1: We propose a novel and compact 4D representation, PackUV, for volumetric videos that packs 3D Gaussian attributes into a sequence of 2D UV atlases (yellow, top right). PackUV is readily compatible with existing video coding infrastructure (e.g., can be coded with HEVC, FFV1). We also propose PackUV-GS, a method to directly fit Gaussian attributes from multi-view RGB videos into structured PackUV (blue, top left) via optical flow-guided keyframing and Gaussian labeling to fit arbitrary length sequences with temporal consistency even in the presence of large motions and disocclusions. The fitted scene can be rendered back to streamable volumetric video from any viewpoint (red, bottom). We also propose PackUV-2B, the largest 4D multi-view dataset containing 2B frames captured with over 50 synchronized cameras to provide 360 coverage.
1 Introduction
Volumetric videos are a form of immersive media that capture scenes in three dimensions and across time (4D), making them viewable from any perspective. They promise numerous applications in AR/VR, entertainment, sports, as well as in applications requiring 4D understanding, for instance, in robotics [98, 68, 23, 20]. Unsurprisingly, creating volumetric videos from multiple camera views has been a long-standing challenge in computer vision and graphics [49, 77, 4, 13, 69].
Common approaches for volumetric video reconstruction rely on explicit representations, such as point clouds [66, 73, 27] meshes [56], multi-plane images [71, 78, 12, 82], or multi-sphere images [3, 6]. However, these methods are limited in their ability to render complex scenes and are highly memory intensive. Meanwhile, radiance fields [50, 88] and 3D Gaussian Splatting  [32] have emerged as leading representations for 3D reconstruction and volumetric video [72, 85, 96, 97, 11, 35]. Despite their success, these methods [85, 41, 90, 97, 96, 11] struggle to operate on videos longer than a few seconds. Streaming approaches [72, 46, 91, 35] address this via online fitting but struggle to retain long-duration temporal consistency, capture large motions, and handle disocclusions (e.g., when a new object enters the scene). Furthermore, volumetric videos produced by these methods cannot be seamlessly shared due to their large size and required bespoke compression methods [15] that are incompatible with existing multimedia infrastructure.
To address these challenges, we introduce PackUV, a novel 4D Gaussian representation that packs 3D Gaussian attributes into a sequence of structured, multi-layered 2D UV maps. These UV map layers are further compacted into a single progressive atlas (Figure 1), enabling efficient storage. We also present PackUV-GS, a fitting method that generates temporally-consistent volumetric videos from multi-view input. Unlike previous 3DGS-to-2D approaches that rely on lossy post-hoc UV unwrapping [65, 86], PackUV-GS directly fits Gaussian parameters into the UV domain. To maintain temporal coherence under large motions and disocclusions, an optical-flow–guided module identifies dynamic Gaussians, enforces flow-based keyframing, and selectively freezes gradients in static regions to ensure stable optimization. This design supports high-quality reconstruction for sequences of arbitrary duration while maintaining quality. The resulting image-native representation replaces point-centric storage with a sequence of ordered, multi-scale 2D UV atlases, enabling efficient streaming and storage via standard video coding methods (e.g., HEVC, FFV1). To the best of our knowledge, this is the first unified representation that applies conventional video coding directly to all 3DGS attributes, with no quality loss during the conversion, to bridge 4D Gaussian representations and the existing video infrastructure.
We evaluate the performance of our method and justify design choices on a variety of existing datasets. However, existing datasets [30, 40, 94, 92, 61] are largely restricted to frontal cameras and exhibit limited motions and disocclusions. To better showcase the abilities of our representation and compare it with existing work, we captured the largest long-duration multi-view video dataset, PackUV-2B. PackUV-2B features real-world dynamic scenes with more than 50 synchronized cameras, providing 360 coverage in both controlled studio and uncontrolled in-the-wild settings. In total, PackUV-2B contains 100 sequences with more than 2B (billion) frames featuring a diverse range of scenarios including human-human, human-object, and human-robot interactions. Extensive experiments show that our method outperforms all baselines across standard metrics and can model much longer sequences (up to 30 minutes) while maintaining consistent quality. To summarize, our contributions include:
  • PackUV, a new volumetric video representation that packs 3D Gaussian attributes into a sequence of UV atlases for efficient streaming and storage, making it readily compatible with existing video coding infrastructure.
  • PackUV-GS, an efficient method to fit PackUV directly from multiview videos using optical-flow-based keyframing and Gaussian labeling to handle large motions, disocclusions, and temporal consistency.
  • PackUV-2B, the largest multi-view 4D dataset with 2B frames, large motions, and disocclusions. It provides 360 coverage from 50+ synchronized cameras.
2 Related Work
In this brief related work, we focus on methods and representations for reconstructing volumetric videos.
4D Volumetric Video.
Volumetric video fitting is a long-standing problem, with works originally focusing on using multi-view images [7, 9, 59, 101], multi-plane images [71, 78, 12, 82], light fields [31, 16, 26, 37] as well as explicit representations like point clouds [66, 73, 27] meshes [56], voxels [17, 18, 53, 55], or multi-sphere images [3, 6]. More recently, radiance fields [50, 5, 10], in particular 3D Gaussian splatting [32], have emerged as the de facto method for static novel view synthesis and reconstruction. Building on 3DGS, a significant amount of work on dynamic scenes has subsequently emerged [28, 34, 43, 45, 97, 85, 87, 95, 19, 93, 11, 25, 70, 90, 39, 62, 14, 2, 29, 33, 42, 58, 81, 60, 80, 40, 67, 24, 8, 79]. Deformable3DGS [97], 4DGS [85], and Grid4D [89] define a deformation field mapping canonical Gaussian primitives to specific time steps. RealTime4DGS [95] and 4D-Rotor-GS [19] introduce 4D Gaussian primitives, improving flexibility for a variety of dynamic scenes. Despite the progress, these methods are limited to short sequences (few seconds), memory intensive, temporally inconsistent, or cannot handle disocclusions (see Section 6).
Recent methods like LongVolCap [92] make tremendous progress by leveraging a hierarchical temporal 4D Gaussian representation to compactly model long-horizon scenes, but still struggle to fit arbitrary durations due to the growth of Gaussians. To address the memory cost of offline 3DGS fitting, 3DGStream [72] and ATGS [11] propose online training and model dynamics of 3D Gaussians per-frame with a neural transformation cache. However, the per-frame storage cost of 3D Gaussians remains large, making it infeasible to transmit and play in a streaming manner like conventional videos. Ex4DGS [36] addresses memory overhead by separating Gaussians into linearly moving ‘static’ and fully dynamic Gaussians. In addition to static-dynamic Gaussian separation, GIFstream [38] and Motion Layering [15] improve dynamic modeling through time-dependent feature streams. However, these methods still struggle to model large motions and disocclusions.
Volumetric Video Representations.
Representing and storing volumetric videos is significantly more expensive than 2D images or videos due to the additional spatial and temporal dimensions. Several approaches have been proposed, either by pruning unused Gaussians, cleverly compressing Gaussian attributes, or using learning [35, 52, 1, 21, 22, 44, 54, 57]. Recently, representations for ‘flattening’ 3D Gaussians into 2D form have been receiving attention, for instance, SOG [51] and UV projection [86, 65]. However, all these methods are limited to static 3D scenes. For 4D volumetric videos a naive sequence of static representations would still be too large.
Recognizing this, some recent works [15, 83] transfer ideas from 2D video coding to volumetric videos. However, due to the unstructured nature of regular 3DGS attributes, these methods have relatively large compression losses, or heavy computational requirements. Structured UV projection methods [65, 86] are promising since they can then be combined with existing image or video coding methods. However, projecting an already-optimized 3DGS into a UV map is lossy and computationally redundant.
Our PackUV representation, together with our native PackUV-GS fitting method, overcomes these limitations by directly producing a sequence of UV atlases that are fully compatible with existing lossless (and lossy) video coding methods (e.g., HEVC, AVC, FFV1). In addition, our method can handle arbitrary length sequences while preserving temporal consistency under large motions and disocclusions.
Figure 2: (Top) Three UV-map organization strategies: (a) naïvely stacking UV layers (deep layers become more and more sparse); (b) a geometric-progression UV pyramid (more uniform sparsity with less storage); (c) PackUV, which packs all pyramid layers into a single UV atlas for efficient, codec-friendly processing. (Bottom) We propose PackUV-GS, a new representation based on 3DGS with a discrete spatial distribution constraint via UV fitting. It uses multiple-layer UV images to store the Gaussian attributes during 3DGS fitting. To constrain the 3D Gaussians located on the discrete rays, we propose a UV-based Adaptive Density Control. We also use a stream-based training schema based on keyframes (image with yellow border).
3 Preliminaries
We first provide a brief background on 3D Gaussian Splatting [32] and UVGS [65] as our method builds upon them.
3D Gaussian Splatting ((3DGS)) [32]:
3DGS is a representation of 3D shape and appearance consisting of a set of Gaussian primitives with a position $\mu\in\mathbb{R}^{3}$, and covariance $\Sigma$. Additionally, each 3D Gaussian primitive explicitly encodes view-dependent appearance via spherical harmonics (SH) coefficients $c$ and an opacity value $o\in\mathbb{R}$. These attributes are optimized by minimizing the loss between the rendered and reference images. To render a viewpoint, the Gaussians are projected as 2D splats and combined with $\alpha$-blending using a tile-based rasterizer.
UV-based Point-to-Image Transformation.
The original 3DGS representation consists of an unstructured set of permutation-invariant primitives. This unstructured nature poses challenges for downstream tasks, particularly when dealing with thousands or millions of Gaussians.
UVGS [65] addresses this by introducing a structured UV-based point-to-image transformation that reformulates 3D Gaussians from an unordered set into a spatially organized UV image via spherical projection. Each Gaussian $g_{i}$ centered at $\mu_{i}=(x_{i},y_{i},z_{i})$ is transformed into spherical coordinates $(\rho_{i},\theta_{i},\phi_{i})$, where
$\rho_{i}=\sqrt{x_{i}^{2}+y_{i}^{2}+z_{i}^{2}}$, $\theta_{i}=\tan^{-1}(y_{i},x_{i})$, and
$\phi_{i}=\cos^{-1}\left(\frac{z_{i}}{\rho_{i}}\right)$. The azimuthal angle $\theta_{i}$ and polar angle $\phi_{i}$ are normalized to discrete UV coordinates in a map of size $M\times N$:
$u_{i}=\left\lfloor\frac{\pi+\theta_{i}}{2\pi}\times M\right\rfloor,\quad v_{i}=\left\lfloor\frac{\phi_{i}}{\pi}\times N\right\rfloor.$ (1)
Since multiple Gaussians may project to the same UV coordinate, UVGS uses multiple layers ($K$ layers) to store top primitives, ordering them by opacity $o$ within each pixel. The final UV mapping is $f:\text{UVCoords}\times\text{LayerIdx}\to\mathbb{R}^{D}$, where $D$ encompasses all Gaussian attributes: $f(u,v,k)=\{\rho,~r,~s,~o,~c\}\in\mathbb{R}^{D}.$
The transformed UVGS representation introduces spatial coherence and resolves the permutation invariance problem. Interestingly, due to opacity-based sorting during the Gaussian mapping process, UVGS can effectively recover the surface-level Gaussians. The capability to represent a set of 3D Gaussian primitives in 2D while preserving surface-level details is of particular importance to our work.
However, it should be noted that this post-optimization UVGS mapping is highly lossy, since it projects only the Gaussian centers (mean positions) onto the UV space. As a result, it performs well for simple 3D objects but applying this post-hoc transformation to pretrained 4D Gaussian sequences often degrades visual quality, causing missing details and temporal inconsistencies (see supplementary).
4 Method
Our goal is to retain the structural benefits of UV mapping, preserve 3DGS’s strong reconstruction quality, and capture 4D dynamic scenes. To achieve this, we first propose PackUV– a novel representation that combines a UV mapping strategy with progressive downscaling to represent 4D volumetric video. Second, we propose a novel fitting method, PackUV-GS (Section 4.2) that uses optical flow based keyframing and Gaussian labeling to ensure smooth and lossless temporally-consistent reconstruction over long-horizon videos. Our approach ensures efficient representation, accurate reconstruction, and scalability to arbitrary length in complex dynamic environments.
4.1 PackUV Atlas
Pyramid UV Mapping:
While direct UV optimization produces structured Gaussian maps, storing all $K$ layers at uniform resolution $M\times N$ incurs significant memory overhead—particularly problematic for high-resolution dynamic sequences. However, we made an important observation: after sorting by opacity, deeper layers (higher $K$) contain progressively fewer visible Gaussians due to occlusion and opacity-based sorting across datasets (more details in the supplementary). This means that not all layers are equally important, and we can therefore adopt a progressive, pyramid-like representation (see Figure 2, top). Instead of storing all $K$ layers at base resolution $M_{0}\times N_{0}$, we apply geometric downsampling that alternates dimension reduction:
$(M_{k},N_{k})=\begin{cases}(M_{0},N_{0}),&k=0\\ (M_{k-1},N_{k-1}/2),&k\text{ odd}\\ (M_{k-1}/2,N_{k-1}),&k\text{ even}\end{cases}.\vskip-5.69054pt$
As shown in Figure 2 (b), this pattern, $M_{0}\times N_{0}$, $M_{0}\times N_{0}/2$, $M_{0}/2\times N_{0}/2$, $M_{0}/2\times N_{0}/4$, …}, reflects increasing sparsity in deeper layers.
UV Atlas Layout
To maximize compactness, we pack the $K$ progressive layers into a single texture atlas $\mathcal{A}$ via recursive subdivision in layout that resembles a quadtree [99] (see Figure 2 (c)):
  • Layer 0: Occupies right region at full resolution $M_{0}\times N_{0}$.
  • Layers 1–2: $L_{1}$ (rotated $90^{\circ}$ CCW) and $L_{2}$ (horizontal) subdivide the left region.
  • Layers 3+: Continue recursive packing rightward of $L_{2}$, alternating orientation (odd layers rotated, even horizontal) at progressively finer resolutions.
This packing technique achieves 88.5% efficiency (utilized pixels / total atlas pixels) while significantly outperforming grid ($\sim$60%) or pyramid ($\sim$75%) layouts. The atlas dimensions we use are:
$W_{\mathcal{A}}=N_{0}+\sum_{k=1}^{K-1}N_{k},\quad H_{\mathcal{A}}=\max_{k}M_{k}.$
Finally, to represent volumetric video, we assemble a continuous sequence of such UV atlases, each one representing 1 video frame. PackUV is seamlessly integrated into our training pipeline: during optimization, UV maps are maintained at their respective progressive resolutions, and upon convergence, layers are packed into the atlas for efficient storage and streaming (Section 4.2).
4.2 PackUV-GS Fitting
The goal here is to efficiently fit PackUV directly from multiview videos using optical-flow-based keyframing and Gaussian labeling to handle large motions, disocclusions, and maintain temporal consistency.
Fitting UV Maps Directly.
Instead of first fitting 3DGS and then projecting to the UV space [65], we optimize the scene Gaussians directly within UV space with fixed spatial resolution and predetermined layer count (K). Let $U\in\mathbb{R}^{M\times N\times K\times D}$ denote the UV maps, where $M\times N$ defines the UV grid resolution, $K$ is the number of layers per pixel, and $D$ encodes all Gaussian attributes:
$U[u_{i},v_{i},k]=g_{i}=\{\rho_{i},~r_{i},~s_{i},~o_{i},~c_{i}\}\in\mathbb{R}^{D}.$ (2)
This direct optimization not only preserves the structural benefits for downstream tasks but also enforces Gaussian sparsity through the discrete UV grid structure.
4.2.1 Video Keyframing
To efficiently fit given synchronized multi-view videos with $T$ frames, represented as a set $\{V_{n}\}_{n=1}^{N}$, where N is the total number of views. We divide each video into s set of $m$ temporal segments. To do so, for each frame $t$, we compute the optical-flow magnitude $M(t)$ on one video, select the top (m1)(m-1) magnitude peaks with a minimum separation $\theta$, and use the first frame of every segment as a keyframe. These keyframes define the segment boundaries. For each keyframe $F^{K}_{i}$, the PackUV Gaussians are initialized from the previous keyframe which preserves temporal and spatial consistency. The frames between keyframes are treated as transition frames. Each transition frame $F^{t}$ is initialized from the preceding frame and refined with a few training iterations compared to $F^{K}$:
$\mathcal{G}(K)\leftarrow\text{Update}\!\left(\mathcal{G}(K-1)\right),\mathcal{G}(t)\leftarrow\text{Update}\!\left(\mathcal{G}(t-1)\right).$
This staged, stream-based strategy enables efficient reconstruction of high-fidelity dynamic scenes while allowing us to parallelize the fitting process. Frames exhibiting high drift, occlusions/disocclusions, or appearance breaks are promoted to keyframes. This keyframing technique helps us handle arbitrary length sequences, large motions, and disocclusions without quality degradation over time. More details are given in the supplementary.
4.2.2 Gaussian Labeling
On top of sequential keyframing pipeline for dynamic Gaussian splatting, we also use optical flow to isolate dynamic regions and freezes static Gaussians during optimization. Flow is computed per camera via RAFT [74], and a CUDA-accelerated covariance-aware projection robustly determines which 3D Gaussians overlap with flow-detected dynamics. The approach improves temporal stability and training efficiency on multi-sequence videos while preserving static backgrounds.
Figure 3: PackUV-GS vs. baselines for large motion and disocclusion handling. The proposed keyframing and Gaussian labeling strategy effectively manages complex scenarios, such as new objects or people entering a room and dispersing. Zoom to view better.
Optical Flow and Binary Motion Masks.
For each camera view $c$, we estimate forward optical flow $\mathbf{F}^{c}_{(t-1)\rightarrow t}$ between consecutive frames $(I^{c}_{t-1},I^{c}_{t})$ using RAFT [74]. We form a binary motion mask by thresholding the flow magnitude and dilating to include local context:
$M_{t}^{c}(\mathbf{p})=\begin{cases}1,&\|\mathbf{F}^{c}_{t-1\rightarrow t}(\mathbf{p})\|_{2}>\tau,\\[4.0pt] 0,&\text{otherwise}.\end{cases}\vskip-5.69054pt$
$\qquad M^{c}_{t}\leftarrow\mathrm{dilate}\bigl(M^{c}_{t};\,r\bigr)\,.$
$\tau$ is the flow magnitude threshold and $r$ the dilation radius.
Covariance-Aware Gaussian Masking:
Each Gaussian $g_{i}$ has mean $\bm{\mu}_{i}\in\mathbb{R}^{3}$, diagonal scales $\mathbf{s}_{i}\in\mathbb{R}^{3}$, and rotation $\mathbf{q}_{i}$ (unit quaternion). Its 3D covariance is represented using $\mathbf{R}(\cdot)$ is the rotation matrix and $\mathbf{S}(\mathbf{s})=\mathrm{diag}(\mathbf{s})$ [32].
Let $\mathbf{T}_{c}$ denote the camera $c$’s view transformation matrix and $\mathbf{J}_{c}(\cdot)$ its $2\times 3$ Jacobian evaluated at the camera-space mean. Following EWA splatting [102], the 2D covariance of $g_{i}$ in image space is
$\Sigma^{2D}_{i,c}\;=\;\mathbf{J}_{c}\,\Sigma^{3D}_{i,\mathrm{cam}}\,\mathbf{J}_{c}^{\top}\,,\qquad\Sigma^{3D}_{i,\mathrm{cam}}=\mathbf{T}_{c}\,\Sigma^{3D}_{i}\,\mathbf{T}_{c}^{\top}\,.$ (3)
We then project the mean to normalized device coordinates (NDC), obtain pixel coordinates $\mathbf{m}_{i,c}\in\mathbb{R}^{2}$, and test overlap with the motion mask using the Mahalanobis metric [47]. A pixel $\mathbf{p}$ is inside the ellipse if
$d^{2}\!(\mathbf{p};\,\mathbf{m}_{i,c},\,\Sigma^{2D}_{i,c})\,=\,(\mathbf{p}-\mathbf{m}_{i,c})^{\top}\bigl(\Sigma^{2D}_{i,c}\bigr)^{-1}(\mathbf{p}-\mathbf{m}_{i,c})\;\leq\;9\,.$
A Gaussian is marked dynamic for camera $c$ if any pixel within a radius derived from the ellipse’s largest eigenvalue satisfies $M^{c}_{t}(\mathbf{p})=1$:
$D_{i,c}\;=\;\bigvee_{\mathbf{p}\in\mathcal{E}_{i,c}}M^{c}_{t}(\mathbf{p})\,,\quad\mathcal{E}_{i,c}=\{\mathbf{p}\,|\,d^{2}\!(\mathbf{p})\leq 9\}\,.$ (4)
The final dynamic mask across cameras is an OR aggregation:
$D_{i}\;=\;\bigvee_{c\in\mathcal{C}}D_{i,c}\,.$
To make the computations realtime, we implement the above in a custom CUDA kernel that, for each Gaussian, (i) computes $\Sigma^{2D}_{i,c}$ via the projection Jacobian, (ii) estimates a sampling radius from the largest eigenvalue, and (iii) scans pixels within this radius, testing $d^{2}\!\leq 9$ and $M^{c}_{t}(\mathbf{p})=1$.
Gradient Freezing.
During backpropagation we zero gradients of all static Gaussians, i.e.,
$\nabla_{\bm{\theta}_{i}}\mathcal{L}\;\leftarrow\;D_{i}\,\nabla_{\bm{\theta}_{i}}\mathcal{L}\,,\qquad D_{i}\in\{0,1\},$
where $\bm{\theta}_{i}$ includes position, scale, rotation, color, and opacity. We additionally reset optimizer momentum for static Gaussians periodically to prevent drift. When densifying, child Gaussians inherit the parent’s dynamic/static label to preserve the dynamic ratio across training.
4.2.3 UV-Based Pruning
To reduce redundancy and keep discretized UV mapping during optimization, we introduce two pruning strategies:
  • Valid UV Projection Pruning. After densification, we recalculate each child Gaussian’s UV coordinates $(u_{\text{scaled}},v_{\text{scaled}})$. Gaussians failing to satisfy equation 1 are pruned. This structural pruning enforces sparsity while aligning Gaussians more closely with scene geometry, improving both efficiency and visual quality.
  • Max-$K$ UV Pruning. To prevent overpopulation at individual UV pixel, we retain only the top $K$ Gaussians per pixel based on opacity. Formally, for each coordinate $(u,v)$ with associated Gaussians $\mathcal{G}_{uv}$:
    $|\mathcal{G}_{uv}|>K\quad\Rightarrow\quad\mathcal{G}_{uv}\leftarrow\text{TopK}(\mathcal{G}_{uv},K)$
These pruning techniques reduce memory overhead, focus learning on surface-relevant regions, and improve reconstruction quality while accelerating optimization.
4.2.4 Objective
Let $\hat{I}^{c}_{t}$ be the rendered image for camera $c$ and ground-truth $I^{c}_{t}$. The photometric objective blends L1 and SSIM:
$\mathcal{L}_{\mathrm{photo}}\;=\;(1-\lambda_{\mathrm{ssim}})\,\lVert\hat{I}^{c}_{t}-I^{c}_{t}\rVert_{1}\;+\;\lambda_{\mathrm{ssim}}\,\bigl(1-\mathrm{SSIM}(\hat{I}^{c}_{t},I^{c}_{t})\bigr)\,.$
To discourage floaters and oversized primitives, we regularize scales and opacities, optionally restricted to dynamic Gaussians ($D_{i}{=}1$):
$\mathcal{L}_{\mathrm{scale}}$ $\;=\;\mathbb{E}_{i}\,\Bigl[\max\{0,\;\max(\mathbf{s}_{i})-s_{\mathrm{max}}\}\Bigr]^{2}\,$, (5)
$\mathcal{L}_{\mathrm{opacity}}$ $\;=\;\mathbb{E}_{i}\,\alpha_{i}\,(1-\alpha_{i})\,.$ (6)
The total loss is
$\mathcal{L}\;=\;\mathcal{L}_{\mathrm{photo}}\;+\;\mathcal{L}_{\mathrm{depth}}\;+\;\lambda_{\mathrm{scale}}\,\mathcal{L}_{\mathrm{scale}}\;+\;\lambda_{\mathrm{opacity}}\,\mathcal{L}_{\mathrm{opacity}}\,.$
Low Precision Optimization.
Unlike prior methods that quantize Gaussian parameters only after training, PackUV-GS performs low-precision optimization (LPO) directly over the attributes $\theta={\mathbf{\mu},\mathbf{s},\mathbf{r},\alpha,\mathbf{c}}$. At each iteration, the renderer consumes a uniformly quantized $K$-bit proxy $\tilde{\theta}$, while gradients flow through a straight-through estimator and FP32 master weights. The training loss (L1+SSIM with a scheduled regularizer) and optimizer (Sparse Adam [48]) remain unchanged. LPO compensates quantization error, preserving both PSNR and training speed. We use 8-bit precision for $\mathbf{s},\mathbf{r},\alpha,\mathbf{c}$} and 16-bit for $\mathbf{x}$}, later split into two 8-bit channels for storage. This 8-bit–per-channel layout is compatible with standard video codecs (e.g., HEVC, AVC, FFV1), enabling both lossless and lossy compression.
Table 1: Dataset Comparisons. We compare our newly captured dataset PackUV-2B with existing multi-view datasets across sequence count, total frames, camera setup, resolution, maximum FPS, scenario type, and view range. PackUV-2B contains 100 diverse sequences totaling over 2B (billion) high-quality frames, recorded with more than 50 cameras at $1920\times 1200$ resolution. The capture system supports up to 90 FPS, providing high temporal fidelity.
Dataset Sequence Frames Camera Resolution Max FPS Scenario View Range
D-NeRF [61] 8 $800\times 800$ Simulator 360
CMU Panoptic [30] 65 $\sim 15$M 31 (+ 480) $1920\times 1080$ 30 Real-World 360
N3DV [40] 6 $\sim 38$K 21 $2028\times 2704$ 30 Real-World face forward
NeRF-DS [94] 8 $\sim 8$K 2 $480\times 270$ Real-World face forward
SelfCap [92] 3 $\sim 1.5$M 22 $3840\times 2160$ 60 Real-World face forward
PackUV-2B 100 $\sim$ 2B+ 55–88 $1920\times 1200$ 90 Real-World 360
5 PackUV-2B Dataset
Existing multi-view datasets [30, 40, 94, 92, 61] mostly use front-facing cameras and offer limited diversity, making them insufficient for evaluating volumetric video methods under fast motion, large deformation, and disocclusion. We introduce PackUV-2B, a real-world, long-horizon multi-view dataset comprising 100 dynamic sequences and over 2Bframes, captured with 55-88 synchronized cameras. The data spans studio and in-the-wild settings, including human–human, human–object, and robot–object interactions. Sequences average 10 minutes and reach up to 30 minutes. To form a comprehensive benchmark, PackUV-2B includes activities with broad variation in motion speed (from slow pouring water to fast basketball and dance), motion scale (from table-top manipulation to pickleball), and object properties (rigid, articulated, reflective, transparent). To our knowledge, PackUV-2B surpasses existing datasets in sequence length, camera count, motion complexity, dynamic diversity, and data volume. We present detailed comparisons between PackUV-2B and existing datasets in Table 1. We plan to release PackUV-2B publicly, with the aim of establishing it as a new benchmark for evaluating general-purpose, long-horizon dynamic reconstruction. Additional details about PackUV-2B can be found in the supplementary material.
6 Experiments
Figure 4: Optical flow. To assess long-term temporal stability, we compute optical flow between consecutive timestamps.
In this section, we conduct comprehensive experiments on three different datasets to verify the effectiveness of our method in generating long volumetric videos with large motion and disocclusion while being compatible with the existing video coding infrastructure. Section 6.1 discusses the qualitative and quantitative evaluations of our proposed method and compare it to the baselines. In Section 6.2, we show how PackUV can be losslessly streamed via existing video infrastructure. The ablation study in Section 6.3 shows the effectiveness of each component in our method.
Setup.
We set the PackUV atlas resolution $M_{0}$ and $N_{0}$ to 1024, and the number of UV layers $K$ to 8. $\theta$ for keyframing is set to 30. $\lambda_{\mathrm{scale}}$ and $\lambda_{\mathrm{opacity}}$ were set to 0.0001.
Dataset.
[…]
Baselines.
We compare with several state-of-the-art baseline methods, 3DGStream [72], 4DGS [84], RealTime4DGS [95], Deformable3DGS [97], ATGS [11], Grid4D [89], Ex4DGS [36], and GIFStream [38]. We train 3DGStream on the full sequence length. To mitigate high VRAM consumption in Deformable3DGS, RealTime4DGS, 4DGS, and Grid4D, we split the sequences into segments and train each segment independently. Although ATGS employs a streaming-framework, we observe that training on full sequences results in gradient explosion and adopt the segmented training strategy. We performed all the experiments on RTX 3090 GPU, 256 GB of RAM and 8 CPU cores. We evaluate rendering quality using average PSNR, SSIM, and LPIPS [100] across all timestamps.
Table 2: Quantitative Comparison. We report PSNR, SSIM, LPIPS, and train time (in hours) for a window length of 60 timestamps. We also report the method’s streaming ability and compatibility with the existing video coding infrastructure.
Method PackUV-2B SelfCap [91] N3DV [40] PSNR SSIM LPIPS Train PSNR SSIM LPIPS Train PSNR SSIM LPIPS Train Stream Codec 3DGStream 23.17 0.826 0.33 1.00 19.77 0.769 0.36 1.43 31.17 0.952 0.23 0.31 Full No 4DGS 23.11 0.808 0.35 2.30 19.56 0.745 0.37 3.18 29.81 0.951 0.21 3.27 No No RealTime 21.37 0.790 0.38 4.48 19.46 0.747 0.41 8.07 32.29 0.955 0.21 2.48 No No Deformable 20.07 0.778 0.33 2.04 17.89 0.708 0.38 2.09 26.51 0.935 0.24 0.62 No No ATGS 21.42 0.796 0.36 1.13 15.48 0.664 0.51 1.82 30.99 0.934 0.24 1.97 Pseudo No Grid4D 21.58 0.790 0.37 1.13 17.53 0.701 0.44 1.82 30.87 0.954 0.197 1.97 No No Ex4DGS 20.73 0.789 0.39 0.83 17.62 0.680 0.39 1.23 31.57 0.944 0.23 0.59 No No GIFStream 21.92 0.795 0.39 1.61 19.78 0.745 0.35 2.05 31.10 0.954 0.25 0.42 Pseudo Partial Ours 27.41 0.842 0.28 1.05 22.52 0.783 0.31 1.12 32.81 0.953 0.21 1.37 Full Full
6.1 Qualitative and Quantitative Results
Qualitative results.
Figure 3 shows qualitative comparisons where our method excels at handling large motions and severe disocclusions. Our keyframing and Gaussian-labeling strategy robustly manages complex events, including new objects or people entering and interacting (e.g., people entering a room and dispersing). Additional results appear in the supplementary material. Deformable3DGS, RealTime4DGS, Grid4D, and 4DGS depend on deformation networks, leading to high VRAM usage, poor scalability to long sequences, and difficulty modeling newly emerging objects. Although streaming-based approaches can support longer sequences, they still produce from artifacts. Specifically, 3DGStream struggles with large motions, ATGS suffers from gradient explosion, and GIFStream contains flickering across the sequentially trained segments. In contrast, PackUV-GS maintains higher and more consistent performance across all motion regimes, enabling high-fidelity free-viewpoint rendering with efficient memory usage. As shown in Figure 3,1 and the supplementary material, PackUV-GS produces sharper, more temporally coherent views.
Quantitative results.
Table 2 quantitatively demonstrates that PackUV-GS outperforms all baselines in visual quality metrics across datasets. While 3DGStream, ATGS, and GIFStream support streaming longer sequences, they still yield suboptimal performance. This demonstrates the effectiveness of our proposed representation and method for handling long-duration sequences, while maintaining compatibility with standard multimedia infrastructure.
Long-term temporal consistency.
To assess long-term temporal stability, we compute optical flow between consecutive timestamps from novel viewpoints. For methods that require a segmented training strategy, we compute optical flow for consecutive timestamps between segments. As shown in Figure 4, this demonstrates the inability of deformation based methods to maintain temporal consistency over time. This further highlights the importance of online fitting methods like ours. It is interesting to note that the rendering quality of 3DGStream and ATGS degrades over time whereas PackUV-GS is consistent.
Table 3: We present quantitative ablation study for various components of our method on PSNR, SSIM, and LPIPS.
Method PSNR SSIM LPIPS Method PSNR SSIM LPIPS
Ours $27.41$ $0.84$ $0.28$ w/o Keyframe $20.95$ $0.77$ $0.38$
w/o UV Optim $23.81$ $0.79$ $0.33$ w/o Labeling $25.42$ $0.82$ $0.31$
No Atlas $27.43$ $0.84$ $0.28$ w/o Codec $27.41$ $0.84$ $0.28$
No LPO $27.52$ $0.85$ $0.27$ - - - -
6.2 Video Coding of PackUV
We can encode PackUV using standard 2D video codecs while preserving quality of the 4D scene. Because PackUV-GS maintains temporally consistent UV layouts and only updates per-pixel attributes over time for dynamic regions, the atlas sequence exhibits strong spatial locality and temporal coherence, allowing direct reuse of mature video coding pipelines.
For each frame in PackUV, we group PackUV layers into 8-bit triplets before encoding them with lossless codecs (e.g., FFV1, HuffYUV) via FFmpeg [76]. Each channel is globally normalized using per-channel min/max computed over the entire sequence. To guarantee bit-exact recovery of the original attributes, we transmit a compact sidecar containing the exact normalization parameters. Decoding is thus invertible: it recovers the atlas sequence, which is de-normalized to reproduce the original values.
This design reduces 4D Gaussian scenes to conventional video assets without custom codecs, enabling storage, streaming, and decoding with off‑the‑shelf tooling. In practice, we obtain a perfect reconstruction with zero error in the lossless setting via FFV1. These results confirm that PackUV enables treating 4DGS as standard video content while retaining faithful scene recovery.
6.3 Ablation Study and Discussion
We perform comprehensive ablations on PackUV-2B to validate each component of our method, examining atlas-based packing, Gaussian labeling, low-precision optimization (LPO), video coding, direct UV-space optimization, and video keyframing (Table 3), using PSNR, SSIM, and LPIPS for evaluation. Removing UV initialization and UV pruning (optimizing only via post-hoc UV projection) leads to a clear drop in PSNR and other metrics, confirming the importance of direct UV-space optimization for fine detail preservation. We also assess mapping from uniform UV-space Gaussians to the PackUV atlas. Because deeper UV layers contain few primitives, compressing them into lower-resolution atlases incurs negligible quality loss.
Our ablations also demonstrate that low-precision optimization is effectively lossless and provides a superior alternative to post-training quantization for efficient Gaussian storage. Moreover, LPO enables easy integration with standard video-coding pipelines without requiring specialized post-processing steps such as in Motion Layering, VCubed, or GIFStream. Because our PackUV representation consists entirely of 8-bit images, it can be directly encoded as video and decoded back without reconstruction loss. Finally, we assess the effect of video keyframing. Resetting gradients at keyframes helps maintain spatial and temporal consistency, preventing the gradual quality degradation observed when training without keyframes. Removing keyframing results in a clear decline in average PSNR and other metrics. Additional results are in the supplementary material.
7 Conclusion
We presented PackUV, a unified 4D Gaussian representation that reorganizes volumetric video into structured UV atlases compatible with standard video codecs, and PackUV-GS, an efficient fitting pipeline that maintains long-term temporal consistency under large motions and disocclusions. By directly optimizing all Gaussian attributes in the UV domain, our approach enables high-quality reconstruction for arbitrarily long sequences while improving streaming efficiency. We further introduced PackUV-2B, the largest long-duration multi-view 4D dataset to date, providing challenging benchmarks for future research. Extensive experiments demonstrate that our method outperforms prior work in quality, scalability, and practicality, offering a promising step toward deployable volumetric video in real-world systems.
ACKNOWLEDGEMENTS
This research was supported by ONR DURIP grant N00014-23-1-2804, NSF CAREER award #\#2143576, and an Amazon Cloud Credits award. We also want to thank Emre Arslan for building the first version of the viewer. We also appreciate the time and efforts of all the participants for data collection.
Supplementary Material
8 PackUV-2B Dataset
We captured PackUV-2B, a novel real-world dataset featuring long-horizon, multi-view video sequences. PackUV-2B comprises 100 such sequences, with over 2B (billion) frames in total, and encompasses a diverse array of scenarios, including human-human interaction, human-object interaction, robot-object interaction, among others. The sequences in PackUV-2B average approximately $10$ minutes in length, with some extending up to $30$ minutes. Notably, to establish PackUV-2B as a comprehensive benchmark for evaluating dynamic reconstruction approaches, we have curated sequences of varying difficulty levels, aiming to cover a broad spectrum of real-world settings. For instance, in terms of motion speed, PackUV-2B includes movements ranging from slow actions like ”inching along” to fast-paced activities such as playing basketball, thereby posing challenges for handling motion blur. Regarding motion scale, PackUV-2B captures both small-scale interactions, like table-top object manipulation, and large-scale activities, such as dance, pickle ball, volleyball, etc. Furthermore, PackUV-2B features diverse object categories, including rigid and articulated objects, as well as some reflective and transparent items. To the best of our knowledge, PackUV-2B significantly surpasses existing datasets in its domain concerning sequence length, number of camera views, motion complexity, dynamic diversity, and overall data volume. More details about the captured sequences are presented in Table 6. The Capture system is discussed in Supp. Sec. 14.
Table 4: PSNR comparison with the baselines on N3DV (flame_salmon), DeskGames, and Technicolor datasets.
Dataset Motion L. 4DGS Ex4DGS 3DGStream GIFStream Ours
N3DV 32.55 32.01 32.11 31.67 28.42 33.06
DeskGames 30.89 30.11 29.48 30.51 29.94 32.74
TechniColor 31.77 28.91 31.33 25.48 27.42 31.87
9 Additional Qualitative and Quantitative Results
More qualitative results are shown in figures 9, 11, 10, 12 at the end of supplementary. From the results, it can be seen that our method, PackUV, significantly outperforms other methods in reconstructing large motions and disocclusions while maintaining the overall quality.
We also evaluate on three additional datasets, including one more sequence from N3DV, Technicolor, and the DeskGames dataset from Motion Layering. A summarized comparison is shown in Tab. 4. Our method performs better across datasets. Metrics are collected from Motion Layering except ours and GIFStream for the same experimental setup.
10 Lossy Post-Optimization UV Mapping for Scenes
Fig. 6 show lossy reconstruction via UVGS [65] mapping for a real-world scene. Even when using 48 layers and 1K resolution of UV maps, there are missing Gaussians resulting in clearly visible artifacts in the renderings. This clearly illustrates the importance of UV mapping during optimizing the Gaussians as done in PackUV.
Table 5: Storage comparison with the baselines (30 frames).
Method LG Grid4D Ex4DGS 3DGStream GIFStream Ours
Storage (MB) 950 76 108 204 16 10
Figure 5: (Left) Compression evaluation via different methods. (Right) PSNR consistency over time.
Figure 6: Post-optimization lossy UV mapping fails to capture details of a real-world scene even with 48 layers and 1K resolution.
11 Video Coding and PackUV Storage
PackUV allow easy and lossless encoding of Volumetric videos using standard 2D video codecs while preserving quality of the 4D scene. Because PackUV-GS maintains temporally consistent UV layouts and only updates per-pixel attributes over time for dynamic regions, the atlas sequence exhibits strong spatial locality and temporal coherence, allowing direct reuse of mature video coding pipelines. This solved one of the major drawbacks of streaming based methods like GIFStream, 3DGStream and AT-GS - high storage requirements of the trained models for every timestamp for lossless conversion. Using the FFV1 codec for lossless video conversion is giving us an average storage rate of under 10 MBPS. Tab. 5 compares storage with SOTA 4DGS methods. Our method outperforms all volumetric video streaming baselines and significantly surpasses specialized per-frame compression methods such as LG [21].
We analyze the compression of PackUV atlases via different techniques including quantization of individual 3DGS attributes, using different video and image compression techniques, and mix of both. The results are presented in Fig. 5(left). FFV1 lossless compression with 8-bit LPO achieves the highest compression ratio with only a 0.11 dB PSNR drop.
Another advantage of PackUV is it’s structural arrangement, which can be used for mapping the scene to a latent space to achieve neural compression [65, 86]. Unlike UVGS [65], GaussianAtlas [86], PackUV allows us to represent the entire scene into a single multiscale atlas. This further reduces the complexity of neural networks required to compress it to a latent space. The objective is to compress a single-layer UV atlas representation—which encodes full PackUV atlas to a compact representation that is both easy to store and invertible back to a 4D scene representation. Since the UV format structurally organizes scene attributes in an image-like format, we directly adopt image-based compression pipelines.
11.1 Neural Compression
Neural Compression is a widely used technique for compressing various modalities [64, 65]. We consider a neural network architecture for compressing PackUV inspired from UVGS [65]. We follow the designing by using three lightweight encoder-decoder networks, each responsible for compressing a specific modality:
  • Position Encoder-Decoder: maps spatial information $\sigma_{i}\in\mathbb{R}^{3\times H_{0}\times W_{0}}$
  • Appearance Encoder-Decoder: maps RGB/SH color and opacity $\{c_{i},o_{i}\}\in\mathbb{R}^{4-45\times H_{0}\times W_{0}}$
  • Covariance Encoder-Decoder: maps scale and rotation $\{s_{i},r_{i}\}\in\mathbb{R}^{7\times H_{0}\times W_{0}}$
Each modality is encoded into a 3-channel latent image per frame. Let $L^{t}\in\mathbb{R}^{M\times N\times 3}$ be the per-frame latent representation at time $t$. These latent images are stored along with the shared decoder weights for the entire scene. This approach offers a consistent compression ratio ranging from [8:1] to [128:1] at the cost of minimal degradation (PSNR drop: [1.5] dB to PSNR drop: [4.5] dB). And the decoding process is very efficient due to the light-weight MLP based decoder.
Table 6: PackUV-2B Dataset Layout. We present our newly captured dataset PackUV-2B sequences . We report number of sequences captured, total timestamps (T), total cameras used to capture the sequence, FPS, setting type, and view range. We also report special tags for each dataset describing the type of activity in the captured sequence. PackUV-2B contains 100 diverse sequences totaling over 2B (billion) high-quality frames, recorded with more than 50 cameras at $1920\times 1200$ resolution. The capture system supports up to 90 FPS, providing high temporal fidelity. Tags: RI - Robot Interaction, HI - Human-human Interaction, OI - Object Interaction, SP - Sports, LM - Large Motion, DO - Disocclusion, TR - Transparent/Reflective Objects, EN - Entertainment .
Sequence Num Sequences Num Cameras FPS Setting Tags View Range
Baby Dance 3 88 30 Studio EN, HI, DO 360
Spot 3 88 30, 90 Studio RI, HI, OI, DO 360
Volleyball 3 88 30, 90 Studio SP, HI, LM, DO 360
Kitchen 3 55 30 Non-studio OI 320
Woodwork 4 55 30 Non-studio OI 320
Pickleball 30 55 30 Non-studio SP, HI, LM 300
Meat Shop 11 84 30 Non-studio OI, DO 320
Kuka Robot 2 84 30 Non-studio RI, OI 300
Panda Robot 3 84 30 Non-studio RI, OI 300
Articulation 2 84 30 Studio HI, OI, DO 360
Chair Play 1 86 30 Studio HI, OI, DO, LM 360
Object Placement 1 88 30 Studio HI, OI, DO, LM 360
Dance 01 3 80 30 Studio HI, DO, LM, EN 360
Dance 02 7 84 30, 60, 90 Studio HI, DO, LM, EN 360
Dance 03 3 82 30, 60, 90 Studio DO, LM, EN 360
Dance 04 3 85 30 Studio DO, LM, EN 360
Yoga 3 78 30 Studio LM, SP 360
Tools Play 2 82 30 Studio LM, DO, OI 360
Board Games 4 86 30 Studio HI, OI, LM, DO 360
Photography 4 88 30 Studio HI, OI, LM, DO 360
Conversation 5 84 30 Studio HI, LM, DO 360
PackUV-2B (TOTAL) 100 55–88 30-90 - - $\sim$360
12 PackUV-GS Optimization Details
12.1 Gaussian Labeling:
We use RAFT [75] is used to estimate the optical flow for each set of images. Flow masks are used to label the Gaussians as dynamic or static and only dynamic Gaussians are optimized during the training. We implemented a custom CUDA kernel to expedite this. The process is Explained in Algorithm 1. As a fallback mechanism, we additionally support a simple point-projection strategy that relies solely on the Gaussian centers, without accounting for their full covariance. In the setting, even when the flow-based masking module is disabled, we observe that training the sequence using a keyframing strategy remains stable and temporally coherent. Specifically, the 3D Gaussians for each new keyframe are initialized using the optimized Gaussian parameters from the preceding keyframe. The same is done for the transition frames. This warm-start initialization propagates geometry and appearance across time, effectively enforcing temporal continuity and preserving structural consistency between consecutive segments, despite the absence of explicit motion-aware supervision. We believe, this is possible by fact that PackUV focus learning on surface-relevant regions and most of them remain static during training of the subsequent timestamps.
Algorithm 1 Covariance-Aware Flow Masking with CUDA Acceleration
1:Gaussian parameters $\{\bm{\mu}_{i},\mathbf{s}_{i},\mathbf{q}_{i}\}_{i=1}^{N}$, camera views $\mathcal{C}$, flow masks $\{M^{c}\}_{c\in\mathcal{C}}$
2:Dynamic mask $D_{i}\in\{0,1\}$ for each Gaussian $i$
3:Initialize $D_{i}\leftarrow 0$ for all $i=1,\ldots,N$
4:for each camera $c\in\mathcal{C}$ do
5:   CUDA Kernel (parallel over all Gaussians):
6:   for each Gaussian $i$ in parallel do
7:     // 1. Compute 3D covariance and transform to camera space
8:     $\Sigma^{3D}_{i}\leftarrow\mathbf{R}(\mathbf{q}_{i})\cdot\text{diag}(\mathbf{s}_{i}^{2})\cdot\mathbf{R}(\mathbf{q}_{i})^{\top}$
9:     $\Sigma^{3D}_{i,\text{cam}}\leftarrow\mathbf{T}_{c}\cdot\Sigma^{3D}_{i}\cdot\mathbf{T}_{c}^{\top}$
10:     $\bm{\mu}_{i,\text{cam}}\leftarrow\mathbf{T}_{c}\cdot\bm{\mu}_{i}$
11:     // 2. Project to 2D using EWA splatting
12:     Compute Jacobian: $\mathbf{J}=\begin{bmatrix}f_{x}/z&0&-f_{x}x/z^{2}\\ 0&f_{y}/z&-f_{y}y/z^{2}\end{bmatrix}$ at $\bm{\mu}_{i,\text{cam}}$
13:     $\Sigma^{2D}_{i,c}\leftarrow\mathbf{J}\cdot\Sigma^{3D}_{i,\text{cam}}\cdot\mathbf{J}^{\top}$
14:     $\Sigma^{2D}_{i,c}[0,0]\leftarrow\Sigma^{2D}_{i,c}[0,0]+0.3$ $\triangleright$ Low-pass filter
15:     $\Sigma^{2D}_{i,c}[1,1]\leftarrow\Sigma^{2D}_{i,c}[1,1]+0.3$
16:     $\mathbf{m}_{i,c}\leftarrow\text{NDC2Pixel}(\mathbf{P}_{c}\cdot\bm{\mu}_{i})$ $\triangleright$ Project mean to pixels
17:     // 3. Check covariance validity
18:     $\text{det}\leftarrow\Sigma^{2D}_{i,c}[0,0]\cdot\Sigma^{2D}_{i,c}[1,1]-\Sigma^{2D}_{i,c}[0,1]^{2}$
19:     if $\text{det}\leq 10^{-7}$ or $\text{det}<0$ or $z\leq 0$ then
20:        $D_{i,c}\leftarrow 0$ $\triangleright$ Invalid, mark static
21:        continue
22:     end if
23:     // 4. Compute sampling radius from eigenvalues
24:     $\text{tr}\leftarrow\Sigma^{2D}_{i,c}[0,0]+\Sigma^{2D}_{i,c}[1,1]$
25:     $\lambda_{\max}\leftarrow\frac{\text{tr}+\sqrt{\max(0,\text{tr}^{2}-4\cdot\text{det})}}{2}$
26:     $r\leftarrow\min\left(r_{\max},\left\lceil 3\sqrt{\lambda_{\max}}\right\rceil\right)$ $\triangleright$ 3-sigma, clamped
27:     // 5. Test ellipse overlap with flow mask
28:     $D_{i,c}\leftarrow 0$
29:     for $\mathbf{p}\in\{\mathbf{m}_{i,c}+\bm{\delta}\mid\|\bm{\delta}\|_{\infty}\leq r\}$ do
30:        if $\mathbf{p}$ within image bounds then
31:          $d^{2}\leftarrow(\mathbf{p}-\mathbf{m}_{i,c})^{\top}(\Sigma^{2D}_{i,c})^{-1}(\mathbf{p}-\mathbf{m}_{i,c})$
32:          if $d^{2}\leq 9$ and $M^{c}(\mathbf{p})>0.5$ then
33:            $D_{i,c}\leftarrow 1$
34:            break
35:          end if
36:        end if
37:     end for
38:   end for
39:   // Aggregate on host (OR across cameras)
40:   for each Gaussian $i$ do
41:     $D_{i}\leftarrow D_{i}\lor D_{i,c}$
42:   end for
43:end for
44:return $\{D_{i}\}_{i=1}^{N}$
12.2 Gradient thresholding:
A common challenge in the existing dynamic 3D reconstruction methods is the uncontrolled growth of the number of Gaussian primitives over time. This progressive accumulation leads to excessive memory consumption and frequent out-of-memory (OOM) errors, especially when fitting long video sequences. To mitigate this issue, we introduce a gradient thresholding mechanism that constrains the total number of Gaussians allowed during optimization. The key idea is to retain only those Gaussians that meaningfully contribute to the reconstruction objective while pruning away redundant or inactive ones.
Let the loss function be denoted by $\mathcal{L}$, and consider a Gaussian primitive $g_{i}$. We compute the gradient norm of its contribution to the loss:
$\|\nabla_{g_{i}}\mathcal{L}\|_{2}$
If this norm falls below a predefined threshold $\tau$, the Gaussian is considered non-contributory and is removed from the scene representation. Formally:
$\text{If }\|\nabla_{g_{i}}\mathcal{L}\|_{2}>\tau,\qquad\text{then }g_{i}\text{ is pruned}$
This strategy ensures that only actively contributing Gaussians are retained during optimization, effectively capping the total number of primitives and maintaining computational efficiency. By doing so, we avoid unnecessary memory overhead and preserve high-quality reconstruction over long temporal horizons.
12.3 Video Keyframing
To efficiently fit given synchronized multi-view videos with $T$ frames, represented as a set $\{V_{n}\}_{n=1}^{N}$, where N is the total number of views. We divide each video into s set of $m$ temporal segments. To do so, for each frame $t$, we compute the optical-flow magnitude $M(t)$ on one video, select the top (m1)(m-1) magnitude peaks with a minimum separation $\theta$, and use the first frame of every segment as a keyframe. These keyframes define the segment boundaries. For each keyframe $F^{K}_{i}$, the PackUV Gaussians are initialized from the previous keyframe which preserves temporal and spatial consistency. The frames between keyframes are treated as transition frames. Each transition frame $F^{t}$ is initialized from the preceding frame and refined with a few training iterations compared to $F^{K}$:
$\mathcal{G}(K)\leftarrow\text{Update}\!\left(\mathcal{G}(K-1)\right),\mathcal{G}(t)\leftarrow\text{Update}\!\left(\mathcal{G}(t-1)\right).$
This staged, stream-based strategy enables efficient reconstruction of high-fidelity dynamic scenes while allowing us to parallelize the fitting process. Frames exhibiting high drift, occlusions/disocclusions, or appearance breaks are promoted to keyframes. This keyframing technique helps us handle arbitrary length sequences, large motions, and disocclusions without quality degradation over time.
Through our experiments we observe that keeping keyframes farther apart can cause the quality to degrade over time, thus we keep the keyframe threshold to 30 timestamps.
12.4 Low-Precision Training
Unlike most prior works that employ lossy quantization after optimization for storage efficiency, PackUV-GS utilizes low precision optimization (LPO) for learning the Gaussian attributes $\theta=\{\mathbf{\mu},\mathbf{s},\mathbf{r},\alpha,\mathbf{c}\}$. Our experiments show that LPO effectively compensates for the quantization loss, maintaining both PSNR and training efficiency. At each iteration, the renderer operates on a quantized proxy $\tilde{\theta}$ obtained by a uniform $K$-bit fake quantizer with scale $\Delta$. We use a straight-through estimator (STE) to preserve gradient flow while keeping master weights in FP32. The training objective remains the same, combining an L1+SSIM photometric term with a scheduled regularizers. Backpropagation updates FP32 master parameters with Sparse Adam optimizer [48, 63].
We use 8-bit reduced precision for $\{\mathbf{s},\mathbf{r},\alpha,\mathbf{c}\}$ and 16-bit for $\{\mathbf{x}\}$. During storage, the 16-bit $\{\mathbf{x}\}$ values are split into two 8-bit parts. This 8-bit-per-channel design makes PackUV readily compatible existing video coding infrastructures, enabling the direct application of both lossless or lossy compression methods (e.g., HEVC, AVC, FFV1). We also conduct several experiments to evaluate the loss incurred from reduced-precision training using different levels of bit quantization. From these experiments, we observe that that quantizing the position parameters ($\{\mathbf{x}\}$) cause the largest PSNR drop during training. Consequently, we retain them in FP16, which introduces negligible error during optimization. In contrast, the other attributes ($\{\mathbf{s},\mathbf{r},\alpha,\mathbf{c}\}$) show no noticeable PSNR drop in our experiments, even when quantized to 8-bit integers after appropriate scaling.
13 Viewer
The viewer renders a dynamic 4D Gaussian scene as a time-ordered sequence of Gaussian frames. At time step $t_{k}$, the scene is represented as
$\mathcal{G}(t_{k})=\{g_{i}^{(k)}\}_{i=1}^{N_{k}},$
where each Gaussian $g_{i}^{(k)}$ stores a 3D mean $\bm{\mu}_{i}$, covariance $\bm{\Sigma}_{i}$, opacity $\alpha_{i}$, and appearance parameters (e.g., SH coefficients). A stream manager loads PackUV-encoded frame data sequentially, maintains a small look-ahead cache, and exposes both CPU-side Gaussian data and GPU-ready packed metadata for the current frame.
During playback, the active frame index is advanced according to the target frame rate,
$k(t)=\lfloor ft\rfloor\bmod N,$
where $f$ is the playback FPS and $N$ is the number of frames. When a new frame is selected, its packed representation is uploaded to the renderer and the Gaussian set is updated before drawing.
Rendering follows the standard 3D Gaussian splatting pipeline. A Gaussian in world coordinates is projected into screen space using the current camera pose and intrinsics. For a 3D point $\bm{X}$, projection is given by
$\bm{X}_{c}=\bm{R}\bm{X}+\bm{t},\hskip 28.80008pt\tilde{\bm{x}}=\bm{K}\bm{X}_{c},$
with image coordinates obtained by perspective division. The projected 2D covariance is approximated as
$\bm{\Sigma}_{2D}=\bm{J}\bm{W}\bm{\Sigma}_{3D}\bm{W}^{\top}\bm{J}^{\top},$
where $\bm{W}$ is the local world-to-camera transform and $\bm{J}$ is the Jacobian of the projection function. The resulting ellipse defines the screen-space footprint of each splat.
To support correct transparency, Gaussians are depth-sorted with respect to the current camera and composited back-to-front:
$\bm{C}=\sum_{i=1}^{M}T_{i}\alpha_{i}\bm{c}_{i},\hskip 28.80008ptT_{i}=\prod_{j<i}(1-\alpha_{j}),$
where $\bm{c}_{i}$ is the color contribution of the $i$-th splat. The viewer additionally supports multiple shading modes, including depth visualization and spherical harmonics shading, making it suitable for interactive inspection of dynamic PackUV-based 4DGS sequences.
Excluding I/O, our viewer achieves >200>\!200 FPS; including I/O, it maintains $\geq\!30$ FPS on a NVIDIA 4090 GPU with a set number of GS in the range of 300K-500K for every scene. The viewer’s rendering performance independent of sequence length due to buffering.
14 CAPTURE System and Camera Synchronization
For capturing the sequences in PackUV-2B, we constructed a dedicated capture studio equipped with $88$ synchronized static cameras, as shown in Figure 7. For the non-studio captures, we also built a wireless version of the similar capture setup. The ultra-large scale of the PackUV-2B dataset raises great challenges to the corresponding data processing steps. To this end, we develop an automatic pipeline to handle camera calibration, color and lighting correction, and automatic synchronization. Notably, the automatic synchronization is efficiently achieved by a carefully designed data structure in Section 14. These cameras are uniformly distributed across the four walls of a rectangular room, enabling the capture of fine-grained details with minimal occlusion. PackUV-2B provides high-resolution frames ($1920\times 1200$) at frame rates of $30$, $60$, or $90$ FPS, selected based on the level of dynamics in each sequence. We plan to make PackUV-2B publicly available, with the goal of establishing it as a new benchmark for evaluating general-purpose, long-horizon dynamic reconstruction.
Figure 7: CAPTURE Studio Layout.
Figure 8: Shows gradient explosion in ATGS training.
CAPTURE
is a designated multi-view capture system with $88$ cameras, suitable for various kinds of capture settings. This section provides an overview of an AVL tree implementation designed specifically for managing data captured by this system. Based on the timecode of each frame data, this implementation achieves automatic synchronization and efficient search. In practice, it is non-trivial and time-consuming to achieve high-accuracy synchronization efficiently when working with such a large number of cameras. We introduces an AVL tree implementation designed specifically for automatically synchronizing, managing and querying frame data based on timecodes. AVL trees are self-balancing binary search trees, ensuring that operations like insertion, deletion, and search can be performed efficiently, typically in $O(logn)$ time, where $n$ is the number of nodes in the tree.
The primary goal of this AVL tree is to transform unstructured raw frame data into an efficient data structure with synchronized frames. This structure allows for:
  1. 1.
    Quick Lookups: Rapidly searching all the frames at some timecode within a specified tolerance (threshold).
  2. 2.
    Data Persistence: Once an AVL tree is built, it can be saved as a binary file to avoid the need to rebuild it from the raw file in the future.
14.1 Implementation
The overview of the implementation can be divided into two steps:
  1. 1.
    Build an AVL tree for each camera;
  2. 2.
    After randomly selecting a reference camera, iterate through all the frames in the reference camera, and search the closest frames from all the other AVL trees as the synchronized frames.
14.2 Build an AVL tree
Following 2, We build an AVL tree for each camera based on the camera information file which stores the correspondence of the frame index $idx_{i}$ and the timecode $t_{i}$. It takes as input each pair of $\{idx_{i},t_{i}\}$, and then inserts it as a node into the AVL tree. The AVL tree’s ensures an insertion logic that the tree remains balanced after each new node is added.
Algorithm 2 Build an AVL Tree from a Camera Information File
1:function BuildAVLTree(filename)
2:   $\text{root node }r\leftarrow\text{NULL}$
3:   Open a camera information file as $\mathcal{F}$
4:   for all $f$ in $\mathcal{F}$ do
5:     Parse $f$ to get a $\{\text{timecode string }t_{i},\text{frame index }idx_{i}$
6:     $r\leftarrow\textsc{InsertIntoAVL}(r,t_{i},idx_{i})$ $\triangleright$ InsertIntoAVL handles node creation and tree balancing
7:   end for
8:   Close $f$
9:   return $r$
10:end function
With a built AVL tree, we can achieve efficient and fast lookups. Since all the data is stored in an AVL tree, which is a type of Binary Search Tree (BST), nodes in a BST are organized such that all nodes in the left subtree of a node have timecodes less than the node’s timecode, and all nodes in the right subtree have timecodes greater. This structure allows for efficient searching. Given a timecode to search, the code iterates through the tree, keeping track of the node it has encountered so far whose timecode is closest to the target timecode. It also considers a threshold to ensure that the ”closest” frame found is “close enough”. If the difference of timecodes between the searched frame and reference frame is larger than a pre-defined threshold, that searched frame will be dropped and viewed as “no synchronized frame”.
Algorithm 3 Find Closest Frame by Timecode
1:function FindClosest(root_node $r$, target_timecode $t$, threshold $\tau$)
2:   closest_node $\hat{r}\leftarrow\text{NULL}$
3:   min_difference $d_{min}\leftarrow\infty$
4:   current_node $r^{\prime}\leftarrow$ root_node $r^{\prime}$
5:   while $r^{\prime}\text{ IS NOT NULL}$ do
6:     difference $d\leftarrow|\text{current\_node.timecode}-\text{target\_timecode}|$
7:     if $d<d_{min}$ then
8:        $d_{min}\leftarrow d$
9:        $\hat{r}\leftarrow r^{\prime}$
10:     end if
11:     if $d=0$ then
12:        break $\triangleright$ Exact match found, cannot be closer
13:     end if
14:     if $t<$ the timecode of $r^{\prime}$ then
15:        $r^{\prime}\leftarrow$ the left child node of $r^{\prime}$
16:     else
17:        $r^{\prime}\leftarrow$ the right child node of $r^{\prime}$
18:     end if
19:   end while
20:   if $\hat{r}\text{ IS NOT NULL}$ and $d_{min}>\tau$ then
21:     return NULL $\triangleright$ No node found within threshold
22:   else
23:     return $\hat{r}$
24:   end if
25:end function
Algorithm 4 Synchronization Workflow
1:procedure SynchronizeAndExtractFrames
2:   Define SyncMetaFile path.
3:   if file at SyncMetaFile exists then
4:     SyncInfo $\leftarrow$ LoadSyncInfo(SyncMetaFile) $\triangleright$ Loaded existing synchronization info.
5:   else
6:     AVLTrees $\leftarrow$ GenerateAVLTrees $\triangleright$ Leverage Algorithm 2
7:     SyncInfo $\leftarrow$ SearchFramesUsingAVLTrees $\triangleright$ Leverage Algorithm 3
8:     SaveSyncInfo(SyncInfo, SyncMetaFile)
9:   end if
10:   ExtractAndSaveImageFramesFromVideos(SyncInfo) $\triangleright$ Extract specific frames directly from videos based on torchcodec.
11:end procedure
14.3 Iteratively Synchronize All the Cameras
After building AVL trees for all the cameras, we prepare an automatic workflow (Algorithm 4) to synchronize all of them. Firstly, we sample a reference camera, either randomly or intentionally. Then, we go through all the frames in the reference camera and look up the closest frame of every AVL tree as the synchronized frame. Finally, we save the synchronization information as a json file for future use. Moreover, based on torchcodec, we achieve efficiently extracting any specific frames directly from raw MP4 video file without the need to extracting all the frames first.
15 Limitations and Future Work
While our method achieves high-quality 4D neural video renderings, some limitations remain that requires further exploration. A primary challenge arises from the inherently unstructured nature of 3D Gaussian representations. Although our approach introduces structure through UV projection, it still requires enforcing a large spatial arrangement to capture the fine details of real-world scenes.
PackUV can sometimes produce dragged Gaussian artifacts when the optical flow is estimated inaccurately, which becomes a bottleneck in the pipeline. Although setting more keyframes and reducing optical flow threshold could fix the problem, improving the flow prediction module or adopting a more robust mapping strategy could lead to better temporal consistency.
Furthermore, while PackUV enables the use of video coding infrastructure by mapping to a single frame, discovering more optimal mappings could further improve storage efficiency. Another promising direction is to make this representation compatible with AR/VR devices, thereby enabling direct 4D streaming for immersive applications.
Figure 9: Baseline comparison on PackUV-2B’s Kitchen sequence.
Figure 10: Shows baseline comparison on SelfCap [91] dataset.
Figure 11: Shows baselines comparison on PackUV-2B’s Baby Dance sequence.
Figure 12: Baselines comparison on PackUV-2B’s SPOT sequence.